OpenStudy (anonymous):

Anyone explain how to factorise: 2x^2-7x-15

OpenStudy (anonymous):

2x^2-7x-15 = (x-5 )(2x +3).

OpenStudy (aroub):

(2x+3)(x-5)

OpenStudy (anonymous):

u need to numbers a and b such that a+b=7/2 ab=-15/2

OpenStudy (anonymous):

(a+b)(c+d)=ac+bc+ad+bd (2x+3)(x-5)=2x^2 +3x -10x -15 =2x^2 7x -15

OpenStudy (anonymous):

Mark O has a good answer but you need to understand how t o work 'back' from his answer in order to factorise fully. Multiply the coefficient of x^2 (which is 2) by the constant -15 giving -30 Find all pairs of factors of -30 then pick the unique pair which add up to -7 (cofficient of x) (-10 and 3) Rewrite the equation with -10x and + 3x in place of -7x 2x^2 -10x + 3x -15 Pair off and bracket... (2x^2 -10x) + (3x -15) now factorise each bracket. 2x(x - 5) +3(x - 5) The factors are (x - 5) [contents of both brackets] (2x + 3) [the remaining elements outside brackets] Answer = (x - 5)(2x + 3)

OpenStudy (anonymous):

Thank you very much "gianfranco", is what I neeeded. Appreciate it :)

OpenStudy (anonymous):

you are more than welcome